### v2

# Supplementary Materials

[…]
main
Supplementary Materials
I Modeling details
I.1 The Model
We model a researcher’s annual productivity as a stochastic process with the following assumptions:
Assumption 1: The number of publications $q_{t}$ in a given year $t$ cannot be negative.
Assumption 2: Initial researcher productivity $q_{0}$ follows an exponential distribution (LABEL:fig:criteriaA).
Assumption 3: The change in productivity from year to year $\delta_{t}=q_{t+1}-q_{t}$ is distributed independently of previous years’ productivity, and follows a Laplace distribution (LABEL:fig:criteriaB-D).
Assumption 4: Researcher productivity exhibits different dynamics in different career stages (LABEL:fig:criteriaE).
Assumption 3 implies that the stochastic process is Markovian. These assumptions do not uniquely define a specific model. For example, the boundary condition around the combinations of Assumptions 1 and 3 could be plausibly modeled with either a reflecting or truncated boundary, and we proceed with a truncated boundary for computational simplicity. We present two possible models that fulfill these assumptions: first, a simplified model, and then, a full model.
Simplified model. We fix the mode of the Laplace distribution globally to an arbitrary constant ($\mu=-1$) and fit the first-year exponential rate parameter using the empirical mean of the data, $\widehat{\lambda}_{0}=4.65$. Then we separate the data into two career stages: years 00 to $4$, representing the “early career” stage, and years $5$ to $20$, representing the remainder of a career. Each career stage has the same location $\mu=-1$, but different scale parameters $\alpha_{1}$ and $\alpha_{2}$, respectively, which are the only free parameters of this simplified model. We then simulate trajectories from the simplified model, and assess whether individual simulated trajectories exhibits the canonical pattern using a simple model selection approach across a family of linear and piece-wise linear regressions. We systematically explore the variance parameter space of this model, and then plot the fraction of trajectories that meet the criteria for being canonical to produce LABEL:fig:simplified_canonical.
We follow Ref. [way2017misleading] in using model selection to classify individual trajectories as canonical. We fit a two-part piecewise linear model to each trajectory for each possible change point between career years 3 and 17, as well as a linear model. We then use the small-sample correction of the AIC (AICc) to determine the best-fitting model. The trajectory is labeled as canonical if the best-fitting model is a piecewise model (that is, the linear model is not the best fit) that additionally fulfills the following criteria: the slope of the first piece is positive, the slope of the second piece is negative, and the magnitude of the slope of the first piece is at least twice the magnitude of the slope of the second piece.
Full model. Rather than assuming only two career stages at a fixed change point, in the full model we allow up to four change points, whose locations are estimated from the data. And we allow both the location and scale parameters of the Laplace distribution to vary with each change point. Lastly, instead of using a constant mode, for mild technical reasons, we estimate a slope parameter $\widehat{\beta}$ for each career stage, where $\widehat{\mu}=\widehat{\beta}q_{t}$, which allows for sharper changes within career stages, which can now be arbitrarily short (Fig. S2).
A career stage $i$ is defined in terms of change points $c_{i},c_{i+1}$, where we construct the data within each career stage $D_{i}=\{(q_{t},q_{t+1})\}$, where $c_{i}\leq t<c_{i+1}$. Each candidate set of career stages is identified by a tuple of change points $(c_{1},c_{2},c_{3})$, where $c_{3}$ or both $c_{2}$ and $c_{3}$ may be omitted. We implicitly assume a change point at the two endpoints for career age 00 and at $t_{\text{max}}$, and we adopt the convention that the career stages include the right endpoints. For instance, the change point set $(2,5)$ would encode career stages 0–2, 3–5, and 6–20. This procedure yields 1,159 possible sets of one, two, or three change points from 1 to 19.
For each career stage $i$, we estimate the model’s parameters using an alternating optimization. First, we estimate a global mode $\widehat{\mu}_{g}$ of the truncated Laplace distribution by identifying the mode of $\mathbb{P}(\delta_{t})$. Second, since the log-likelihood of the Laplace is then both smooth and convex in the scale parameter for each career stage $\alpha_{i}$, we perform maximum likelihood estimation on $\alpha_{i}$ using $\widehat{\mu}_{g}$ as the mode. Third, we parameterize the log-likelihood in terms of a slope $\widehat{\beta}$, rather than location, for each career stage, writing $\mu_{i}(q_{t})=\beta_{i}q_{t}$, and setting $\alpha_{i}=\widehat{\alpha}_{i}$ from the previous step. Finally, we re-estimate $\alpha_{i}$ using $\widehat{\mu}_{i}=\widehat{\beta}_{i}q_{t}$.
We estimate separate parameters for each career stage. A career stage $i$ is defined in terms of change points $c_{i},c_{i+1}$, where we construct the data within each career stage $D_{i}=\{(q_{t},q_{t+1})\}$, where $c_{i}\leq t<c_{i+1}$. In order to accommodate the flexibility of empirical career structures while remaining computationally feasible, we consider every possible set of four or fewer career stages, or equivalently, one, two, or three change points. We select career stages via model selection using the AIC to correct for overparameterization. Thus each candidate set of career stages is identified by a tuple of change points $(c_{1},c_{2},c_{3})$. For each career stage $i$, we estimate the model’s parameters using an alternating optimization.
We estimate a single exponential parameter $\widehat{\lambda}_{0}>0$ for the distribution of $q_{0}$, which is shared across all models and is independent of the choice of change points. To generate a synthetic productivity trajectory, we first generate a choice of initial productivity $q_{0}$ from the estimated model $\mathbb{P}(q_{0}|\lambda_{0})$. Then we set $q_{t+1}=q_{t}+\delta_{t}$ for $c_{i}<t\leq c_{i+1}$ which we draw from the parameterized distribution $\mathbb{P}(\delta_{t}|q_{t},\alpha_{i},\beta_{i})$. The synthetic productivity trajectories simulated from the model are used to perform the comparisons with the empirical data shown in LABEL:fig:fit and LABEL:fig:unfit.
I.2 Validity of the Markovian assumption
The Markov assumption implies that
$P(q_{t+1}=r_{t+1}$ $|q_{t}=r_{t},q_{t-1}=r_{t-1})$ (1)
$=P(q_{t+1}=r_{t+1}$ $|q_{t}=r_{t}).$ (2)
By conditioning on the intermediate value $q_{t}$, we can observe whether $\mathbb{P}(q_{t+1})$ varies by the value in the previous timestep, $q_{t-1}$ (Fig. S1). We find confirmation for two of the results stated in the main paper. First, compared to simulated trajectories, the empirical trajectories have less variance within careers, greater variance across careers, and many more years with zeros (LABEL:fig:unfit). Here we confirm this observation from the point of view of the Markov property: the overall trend of increasing means implies that historically high (resp. low) values lead future values to be high (resp. low), even after conditioning on intermediate timesteps (Fig. S1). In the main paper, we find significantly greater abundance of years with zero papers empirically than our model predicted, even after correcting for missingness in the data (LABEL:fig:unfitB,D). We still see that the non-Markovian behavior is strongest around zero (that is, conditioning on $q_{t}=0$, we see that trajectories where $q_{t-1}=0$ are far more likely to produce zeros in year t+1 than other values of $q_{t-1}$), but nevertheless there is a smaller non-Markovian trend in the other conditional distributions as well, particularly in the top-most quantiles. This provides validation for our explanation for how empirical trajectories reach higher cumulative productivities by year 5 than the simulations would predict (LABEL:fig:unfitC).
Figure S1: Conditional distributions of productivity. Violin plots showing the conditional distribution and box-and-whiskers of quantiles of productivity, given the previous year productivity. Conditional distributions are shown as kernel density estimates, and the boxes display quartiles of the data, with the median as a white dot. The whiskers are the farthest datapoint within 1.5 times the interquartile range from the nearest hinge. Higher quantiles are more productive.
I.3 The distribution of the random walk increments
Figure S2: Distribution of increments given productivity. (A) For two Laplace distributions with the same estimated mode ($\mu=-2$), the means (horizontal lines) of the distribution with higher variance ($\alpha=4$, solid black) are higher than the means of the distribution with lower variance ($\alpha=2$, dashed pink). (B) A similar diagram, except the location is parameterized in terms of $\beta$, where $\mu=\beta q_{t}$. Here $\beta=-1$. The difference in location is emphasized on the plots as thick green lines, where in (A) the line is constant at $\mu=-2$, while in (B) the line has a negative slope of $\beta=-1$.
For an observation $(x,\delta):=(q_{t},q_{t+1}-q_{t})$, the conditional pdf from the Laplace distribution with a fixed mode $\mu$ (Fig. S2A) is:
$f(x,\delta)=\mathbb{I}\{\delta>-x\}(2-e^{-\mu/\alpha})^{-1}(1/\alpha)e^{-|\delta-\mu|/\alpha}$
On the other hand, the conditional pdf from the Laplace distribution with a productivity-dependent mode $\beta$ where $\mu=\beta x$ (Fig. S2B) is given by:
$f(x,\delta)=\mathbb{I}\{\delta>-x\}(2-e^{-\beta x/\alpha})^{-1}(1/\alpha)e^{-|\delta-\beta x|/\alpha}$
We employ the fixed mode $\mu$ for the simplified model, since it is simpler, and the productivity-dependent mode $\beta$ for the full model, since it produces a more qualitatively successful fit of the empirical data. In particular, using the $\mu$ parameterization, the tails of the annual fluctuations are sufficiently heavy that the convergence rate is unrealistically slow for the timescales representing real productivity trajectories, and the $\beta$ parameterization allows for faster convergence.
The estimated $\beta_{i}$ parameters from the full model are less than $1$ in every career stage, even the first stage where average productivity increased annually, meaning that typical annual productivity decreases compared to the prior year’s productivity. In other words, faculty appear to experience a general “drag” in maintaining their productivity, which we interpret as asymmetry between decreasing productivity (requiring only inaction) compared to increasing productivity (requiring an increase in resources and effort). It is higher variance, made asymmetric by the hard boundary at productivity zero, that overcomes this drag and effectively prevents faculty productivity from slumping toward zero.
II Recovering known parameters from synthetic data
Figure S3: Model recovery on synthetic data with known structure. Parameter estimates and simulated productivity trajectories for three qualitatively different specifications of the generative model, where (A) stages are delineated by sharp and linear changes, (B) by less sharp but nonlinear changes, or (C) simulated trajectories that were generated using the parameters estimated from the empirical data, illustrating that in realistic parameter regimes we can reasonably recover the same parameters.
Using synthetic trajectories with known parameters, the full model correctly recovers the relevant parameters generally up to an error of ±$\pm 0.1$ for $\lambda$ and $\alpha$, and ±$\pm 0.01$ for $\beta$ (Fig. S3). To put that in perspective, an error in $\lambda$ of $0.1$ corresponds to a tenth of a publication in a scientist’s first year, and the errors in $\alpha$ and $\beta$ can be interpreted on a similar scale. This scale of error is practically negligible compared to the much larger underlying variability in productivity across individuals.
Trajectories generated from the fitted model closely align with the original synthetic data across a qualitatively diverse range of scenarios with a varying number of change points (Fig. S3). The inferred change points match the change points used to generate the original synthetic data, and we select the correct number of change points in each of these instances, both in situations with a sharp cutoff like in scenario (A), as well as when change points are more subtle as in scenarios (B) and (C).
III DBLP Data
This random walk model can be applied to any dataset composed of time series of individual researcher productivities that range onward from the beginning of a career, e.g., the first year of a permanent research position like a tenure-track job at a PhD-granting institution. Datasets of individual-level productivity derived from bibliographic databases like the Web of Science [clarivate2022wos] are attractive because of their scale, but pose additional complexities due to the need to first disambiguate individuals, then stratify by fields, which can exhibit widely different average productivity levels [piro2013macro], and finally stratify by different roles, e.g., faculty vs. trainees, or researchers employed at institutions with different research intensities.
We avoid such complexities by studying a dataset of known computer science faculty at PhD-granting institutions in the US or Canada [way2017misleading], which we linked to their publications as recorded in DBLP [dblp2016november], a bibliographic database that focuses on computing and which was used to identify the underlying diversity of productivity trajectories [way2017misleading]. The faculty dataset was a complete collection of the 205 department or school-level academic units on the Computing Research Association’s Forsythe List of PhD-granting departments in computing-related disciplines in the US and Canada [clauset2015systematic, cra2012forsythe]. The manual collection process yielded 5,032 faculty in these units, and the dataset was subsetted to the 2,583 whose PhD degree and first assistant professorship appointment were at units in the sample. The DBLP was then joined to the faculty census, using manual name disambiguation as necessary, yielding 2,453 faculty with linked publication records. The further inclusion criteria described in the “Data” section of the main manuscript resulted in the reported 2,085 faculty in our analysis.
Current faculty often begin publishing before the start of their faculty career, but the underlying dynamics of productivity are different in the pre-faculty stage of a career [zhang2022labor]. We focus our analysis here only on productivity during a faculty career, and so we exclude those publications from years prior to that career’s beginning. Although DBLP has reasonably good coverage over most venues in which computer science faculty publish, it is not complete, and faculty working in subfields that are not well-indexed by DBLP, such as interdisciplinary computing, would appear as anomalously unproductive researchers.
In order to correct for rising individual productivity and uneven historical coverage, Way et al. adjusted historical productivities using a linear model estimated from publication data manually extracted from CVs [way2017misleading]. Such a linear adjustment preserves years with zero papers, which may lead to an over-estimate in the number of years with zero publications. To evaluate this possibility, we randomly select eight individuals among those with at least ten years of zero publications, and manually counted the number of years when they have zero publications in both DBLP (108) and on their websites and Google Scholar (82). This tally yields a maximum likelihood binomial estimate that $75.9\%$ (95% CI: $(67.9\%,84.0\%)$) of DBLP-observed zero-publication years are correct. With the most conservative correction ($67.9\%$ of years with zero publications are truly zero), and most conservative estimate for the empirical mean number of zeros ($14.1\%$ at the bottom of the estimated 95% CI), we would still find that $9.5\%$ of years in the full empirical trajectories are years with zero publications. This estimate excludes the 95% CI of the mean number of zeros in 10,000 simulated trajectories $(8.5\%,8.7\%)$, implying that zero-productivity years are likely more common in real productivity trajectories than our model can account for.